Find the standard deviation of the data summarized in the given frequency distribution.

The manager of a bank recorded the amount of time each customer spent waiting in line during peak business hours one Monday. The frequency distribution below summarizes the results. Find the standard deviation. Round your answer to one decimal place.
Waiting time Number of (minutes) customer 0-3 4-7 8-11 12-15 16-19 20-23 No co o

Group of answer choices

7.0 min

4.5 min

4.8 min

5.0 min
Answer #1

Midpoint of classes = (Lower limit + upper limit ) / 2

So,

Midpoint for class 0 - 3 = 0 + 3 / 2 = 1.5

Similarly, midpoint for class 4-7 = 5.5

midpoint for class 8-11 = 9.5

midpoint for class 12-15 = 13.5

midpoint for class 16-19 = 17.5

midpoint for class 20-23 = 21.5

Let midpoint be X and number of customers be f

Mean = \sum X * f / \sumf

= (1.5 * 9 + 5.5 * 16 + 9.5 * 15 + 13.5 * 8 + 17.5 * 0 + 21.5* 2) / 50

= 7.9

Standard deviation = sqrt [ \sum X2 * f - n * mean2 / n-1 ]

= sqrt [ (1.52 * 9 + 5.52 * 16 + 9.52 * 15 + 13.52 * 8 + 17.52 * 0 + 21.52 * 2) - 50 *7.92 / 49 ]

= 4.8